# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ Appendix A Auxiliary results
If $(h_{n},ABh_{n})\rightarrow(h,k)\in\mathcal{H}\oplus\mathcal{H}$, then $(Bh_{n},ABh_{n})\rightarrow(Bh,k)$ by continuity of $B$. Since $A$ is closed, $Bh\in\operatorname{dom}{A}$, that is, $h\in\operatorname{dom}{AB}$, and $k=ABh$. Next, if $(h_{n},Ah_{n}+Bh_{n})\rightarrow(h,k)\in\mathcal{H}\oplus\mathcal{H}$, then $Bh_{n}\rightarrow z$ for some $z\in\mathcal{H}$ by continuity of $B$, and $(h_{n},Ah_{n})\rightarrow(h,k-z)$. Since $A$ is closed, $h\in\operatorname{dom}{A}=\operatorname{dom}{A+B}$ and $Ah=k-z=k-Bh$. Finally, we have $\{(h,Ah):\ h\in\operatorname{dom}{A}\}=\{(A^{-1}k,k):\ k\in\mathcal{H}\}$, and this set is closed if and only if $\{(k,A^{-1}k):\ k\in\mathcal{H}\}$ is closed. ∎
@@
Let $\mathcal{H}$ be a Banach space. If $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ and $\operatorname{dom}{A}=\mathcal{H}$, then $A$ is continuous if and only if $A$ is closed.
@@
It follows by definition of continuity that $A$ is closed for $A\in\mathcal{B}(\mathcal{H})$. For the converse, see [7, Theorem III.12.6]. ∎
Definition A.12.
Let $\mathcal{H},\mathcal{K}$ be a separable Hilbert space and $A:\mathcal{H}\rightarrow\mathcal{K}$ a densely defined linear operator on $\mathcal{H}$. Then we define
$\operatorname{dom}{A^{*}}:=\{k\in K:\ h\mapsto\langle Ah,k\rangle\text{ is a bounded linear functional on }\operatorname{dom}{A}\}.$
As $\operatorname{dom}{A}\subset\mathcal{H}$ dense, if $k\in\mathcal{K}$, there exists by the Riesz representation theorem a $f\in\mathcal{H}$ such that $\langle Ah,k\rangle=\langle h,f\rangle$ for all $h\in\mathcal{H}$. We define $A^{*}:\operatorname{dom}{A^{*}}\rightarrow\mathcal{H}$ by setting $A^{*}k=f$.
@@
Let $\mathcal{H}$ be separable Hilbert. If $A,B:\mathcal{H}\rightarrow\mathcal{H}$ are densely defined, then
@@
Let $k\in\operatorname{dom}{B^{*}A^{*}}$ and $h\in\operatorname{dom}{AB}$. Since $k\in\operatorname{dom}{A^{*}}$ and $Bh\in\operatorname{dom}{A}$, $\langle ABh,k\rangle=\langle Bh,A^{*}k\rangle.$ Since $A^{*}k\in\operatorname{dom}{B^{*}}$ and $h\in\operatorname{dom}{B}$, $\langle Bh,A^{*}k\rangle=\langle h,B^{*}A^{*}k\rangle.$ Thus $\langle ABh,k\rangle=\langle h,B^{*}A^{*}k\rangle$. Hence $h\mapsto\langle ABh,k\rangle$ is bounded and $(AB)^{*}k=B^{*}A^{*}k$, proving $(i)$. If $B^{*}A^{*}\in\mathcal{B}(\mathcal{H})$, then $\operatorname{dom}{(AB)^{*}}\subset\mathcal{H}=\operatorname{dom}{B^{*}A^{*}}$, showing $(ii)$. ∎
Definition A.14.
Let $\mathcal{H}$ be a separable Hilbert. A densely defined operator $A:\mathcal{H}\rightarrow\mathcal{H}$ is said to be symmetric if $\langle Ah,k\rangle=\langle h,Ak\rangle$ for all $h,k\in\operatorname{dom}{A}$. If $A=A^{*}$, then $A$ is said to be self-adjoint.
@@
Lemma A.16.
[7, Proposition X.2.4] Let $H$ be a separable Hilbert space and $A$ a symmetric operator on $\mathcal{H}$.
@@
Let $\mathcal{H}$ be a Hilbert space and $A:\mathcal{H}\rightarrow\mathcal{H}$ have a bounded inverse $A^{-1}\in\mathcal{B}(\mathcal{H})$. Then $\sigma(A)=\sigma(A^{-1})^{-1}$.
Proof.
Let $\lambda\in\mathbb{C}$. Then $A-\lambda^{-1}I=\lambda^{-1}(\lambda A-I)=-\lambda^{-1}A(A^{-1}-\lambda I).$ Thus $A-\lambda^{-1}I$ has a bounded inverse if and only if $A^{-1}-\lambda I$ has a bounded inverse. ∎
Lemma A.18.
Let $\mathcal{H}$ be a Hilbert space and $A:\mathcal{H}\rightarrow\mathcal{H}$ injective. Then the point spectrum $\sigma_{p}(A)\coloneqq\{\lambda:\ Av=\lambda v\text{ for some }v\not=0\}$ satisfies $\sigma_{p}(A)^{-1}\coloneqq\{\lambda^{-1}:\ \lambda\in\sigma_{p}(A)\}=\sigma_{p}(A^{-1})$.
Proof.
Since $A$ is injective, $0\not\in\sigma_{p}(A)$ and $A^{-1}:\operatorname{ran}{A}\rightarrow\mathcal{H}$ is well-defined. Furthermore, $(A-\lambda)v=0$ if and only if $(A^{-1}-\lambda^{-1})v=0$ for $\lambda\in\sigma_{p}(A)$. Thus, $\lambda\in\sigma_{p}(A)$ if and only if $\lambda^{-1}\in\sigma_{p}(A^{-1})$. ∎
Lemma A.19.
Let $\mathcal{H}$ be a separable Hilbert space and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{1}(\mathcal{H})_{\mathbb{R}}$ be nonnegative. If $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}\subset\mathcal{H}$ densely, then the following holds.
@@
  3. (iii)
    If $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, then $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}\in\mathcal{B}(\mathcal{H})$.
@@
By Lemma A.16 part (i), $\mathcal{C}_{1}^{1/2}$ and hence $\mathcal{C}_{1}$ are injective. So (i) holds. By part (ii) of the same lemma, $\mathcal{C}_{1}^{-1/2}$ and $\mathcal{C}_{1}^{-1}$ are bijective and self-adjoint. The inverse of a compact operator on its range in an infinite-dimensional space is unbounded, hence (ii) holds. If $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$, then $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is well-defined, and by Lemma A.10, $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is a closed operator, and hence bounded by Lemma A.11, showing (iii). ∎
Condition $(i)$ of the Feldman–Hajek theorem Theorem 3.1 can be stated equivalently as follows.
@@
Let $\mathcal{H}$ be a Hilbert space and $\mu=N(m_{1},\mathcal{C}_{1})$ and $\nu=(m_{2},\mathcal{C}_{2})$ be Gaussian measures on $\mathcal{H}$ such that $\ker{\mathcal{C}_{1}^{1/2}}=\{0\}=\ker{\mathcal{C}_{2}^{1/2}}$. Then $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$ if and only if $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is a well-defined invertible operator in $\mathcal{B}(\mathcal{H})$.
@@
Suppose that $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Then $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is well-defined and bijective. By Lemma A.16 part $(ii)$, $\mathcal{C}_{1}^{-1/2}$ is self-adjoint, hence closed. By Lemma A.10, $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is closed, and by Lemma A.11, it is bounded. Conversely, if $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}\in\mathcal{B}(H)$ is invertible, then,
$\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ =$=\{\mathcal{C}_{1}^{1/2}h:h\in\mathcal{H}\}=\{\mathcal{C}_{1}^{1/2}\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}h:h\in\mathcal{H}\}$
=$=\{\mathcal{C}_{2}^{1/2}h:h\in\mathcal{H}\}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}.$
Lemma A.21.
Let $\mathcal{H}$ be a seperable Hilbert space and $\mu=N(0,\mathcal{C})$ a Gaussian measure on $\mathcal{H}$. If $X\sim\mu$ and $\mathcal{C}=SS^{*}$ for $S\in L_{2}(\mathcal{H})$. Then $\mathbb{E}\lVert X\rVert^{2}=\lVert S^{*}\rVert_{L_{2}(\mathcal{H})}=\lVert S\rVert_{L_{2}(\mathcal{H})}^{2}.$
Proof.
Let $(e_{i})_{i}$ an ONB of $\mathcal{H}$ and $X=\sum_{i}\langle X,e_{i}\rangle e_{i}$. Then by Tonelli’s theorem, the definition of the covariance operator, the hypothesis that $\mathcal{C}=SS^{*}$, and the invariance of the Hilbert–Schmidt norm under adjoints,
$\mathbb{E}\lVert X\rVert^{2}$ $=\mathbb{E}\sum_{i}\lvert\langle X,e_{i}\rangle\rvert^{2}=\sum_{i}\mathbb{E}\lvert\langle X,e_{i}\rangle\rvert^{2}=\sum_{i}\langle\mathcal{C}e_{i},e_{i}\rangle=\sum_{i}\lVert S^{*}e_{i}\rVert^{2}=\lVert S^{*}\rVert_{L_{2}(\mathcal{H})}^{2}=\lVert S\rVert_{L_{2}(\mathcal{H})}^{2}.$
Lemma A.22.
If $\mathcal{H}_{1},\mathcal{H}_{2}$ are seperable Hilbert spaces, $X\sim\mu=N(\mu,\mathcal{C})$ Gaussian distribution on $\mathcal{H}_{1}$ and $A\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$, then the distribution of $AX$ is $N(Am,ACA^{*})$
Lemma A.23.
Let $\mathcal{H}$ be a Hilbert space, $(e_{i})_{i}$ an orthonormal sequence and $(\delta_{i})_{i}\in\ell^{2}((-1,\infty))$. Then $I+\sum_{i}\delta_{i}e_{i}\otimes e_{i}$ is invertible in $\mathcal{B}(\mathcal{H})$, with inverse $I-\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}$.
Proof.
Note that $\frac{\delta_{i}}{1+\delta_{i}}\leq c\cdot\delta_{i}$, with $c^{-1}=\inf_{i}{1+\delta_{i}}>0$, by the hypothesis that $(\delta_{i})_{i}\in\ell^{2}((-1,\infty))$. Thus, $\frac{\delta_{i}}{1+\delta_{i}}\rightarrow 0$ and $\sum_{i}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}$ is well-defined. For $h\in\mathcal{H}$, we have by direct computation,
$=\sum_{i}\left(1-\frac{\delta_{i}}{1+\delta_{i}}+\delta_{i}-\frac{\delta_{i}^{4}}{1+\delta_{i}}\right)\langle h,e_{i}\rangle e_{i}$
$=\sum_{i}\langle h,e_{i}\rangle e_{i}=h.$
Similarly,
$\left(I-\sum_{i\in\mathbb{N}}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}\right)\left(I+\sum_{j\in\mathbb{N}}\delta_{i}e_{i}\otimes e_{i}\right)h=$ $\left(I-\sum_{i\in\mathbb{N}}\frac{\delta_{i}}{1+\delta_{i}}e_{i}\otimes e_{i}\right)\sum_{j\in\mathbb{N}}\left(1+\delta_{i}\right)\langle h,e_{j}\rangle e_{j}$
$=$ $\sum_{i\in\mathbb{N}}\left(1+\delta_{i}-\frac{\delta_{i}}{1+\delta_{i}}(1+\delta_{i})\right)\langle h,e_{i}\rangle e_{i}.$
@@ B.1 Proofs of Section 3
The statements about extensions of operators follow from Lemma A.13.
By the Feldman–Hajek theorem, Theorem 3.1, $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Thus, by Lemma A.20, $A:=\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is a well-defined bounded and invertible operator, and by Theorem 3.1(iii), $AA^{*}-I$ is Hilbert–Schmidt. That is, there exists a sequence $(\lambda_{i})_{i}\subset\ell^{2}(\mathbb{R})$ and ONB $(w_{i})_{i}$ of $\mathcal{H}$ such that,
@@
As $A$ is invertible in $\mathcal{B}(H)$, so are $A^{*}$ and $AA^{*}$. But $AA^{*}>0$, so $AA^{*}-I>-1$, showing that $\lambda_{i}>-1$ for all $i$, which proves item (i) holds.
@@
Notice that $\frac{-\lambda_{i}}{1+\lambda_{i}}\in\ell^{2}((-1,\infty))$, since $1+\lambda_{i}\rightarrow 1$ and $(\lambda_{i})_{i}\in\ell^{2}((-1,\infty))$. This proves items (ii), (iii) and (iv). ∎
See 3.6
Proof of Lemma 3.6.
By Lemma 3.3 with $\mathcal{C}_{1}\leftarrow\mathcal{C}_{\textup{pr}}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, there exists an eigenvalue sequence $(\lambda_{i})_{i\in\mathbb{N}}\subset\ell^{2}((-1,\infty))$ and ONBs $(w_{i})_{i\in\mathbb{N}}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ and items (i), (ii), (iii) and (iv) of Lemma 3.3 hold. By (3c), it holds on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$,
@@
$\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}-I=\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}-\mathcal{C}_{\textup{pos}}^{-1})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(-H)\mathcal{C}_{\textup{pos}}^{1/2}.$

### v3

@@ A.2 Unbounded operators
If $(h_{n},ABh_{n})\rightarrow(h,k)\in\mathcal{H}\oplus\mathcal{H}$, then $(Bh_{n},ABh_{n})\rightarrow(Bh,k)$ by continuity of $B$. Since $A$ is closed, $Bh\in\operatorname{dom}{A}$, that is, $h\in\operatorname{dom}{AB}$, and $k=ABh$. This shows item (i). Next, if $(h_{n},Ah_{n}+Bh_{n})\rightarrow(h,k)\in\mathcal{H}\oplus\mathcal{H}$, then $Bh_{n}\rightarrow z$ for some $z\in\mathcal{H}$ by continuity of $B$, and $(h_{n},Ah_{n})\rightarrow(h,k-z)$. Since $A$ is closed, $h\in\operatorname{dom}{A}=\operatorname{dom}{A+B}$ and $Ah=k-z=k-Bh$. This shows item (ii). Finally, if $A$ is also injective, then we have $\{(h,Ah):\ h\in\operatorname{dom}{A}\}=\{(A^{-1}k,k):\ k\in\operatorname{ran}{A}\}$, and this set is closed if and only if the set $\{(k,A^{-1}k):\ k\in\operatorname{ran}{A}\}=\{(k,A^{-1}k):\ k\in\operatorname{dom}{A^{-1}}\}$ is closed. This shows item (iii). ∎
@@
Let $\mathcal{H}$ be a Banach space. If $A:\operatorname{dom}{A}\subset\mathcal{H}\rightarrow\mathcal{H}$ satisfies $\operatorname{dom}{A}=\mathcal{H}$, then $A$ is continuous if and only if $A$ is closed.
@@
It follows by definition of continuity that $A$ is closed for $A\in\mathcal{B}(\mathcal{H})$. For the converse, see [8, Theorem III.12.6]. ∎
Definition A.18 ([8, Definition X.1.5]).
Let $\mathcal{H},\mathcal{K}$ be separable Hilbert spaces and $A:\mathcal{H}\rightarrow\mathcal{K}$ be a densely defined linear operator on $\mathcal{H}$. Then we define
$\operatorname{dom}{A^{*}}\coloneqq\{k\in\mathcal{K}:\ h\mapsto\langle Ah,k\rangle\text{ is a bounded linear functional on }\operatorname{dom}{A}\}.$
As $\operatorname{dom}{A}\subset\mathcal{H}$ is dense, if $k\in\mathcal{K}$, there exists by the Riesz representation theorem some $f\in\mathcal{H}$ such that $\langle Ah,k\rangle=\langle h,f\rangle$ for all $h\in\mathcal{H}$. We define $A^{*}:\operatorname{dom}{A^{*}}\rightarrow\mathcal{H}$ by setting $A^{*}k=f$.
@@
Let $\mathcal{H}$ be a separable Hilbert space. If $A,B:\mathcal{H}\rightarrow\mathcal{H}$ are densely defined, then
@@
Let $k\in\operatorname{dom}{B^{*}A^{*}}$ and $h\in\operatorname{dom}{AB}$. Since $k\in\operatorname{dom}{A^{*}}$ and $Bh\in\operatorname{dom}{A}$, $\langle ABh,k\rangle=\langle Bh,A^{*}k\rangle.$ Since $A^{*}k\in\operatorname{dom}{B^{*}}$ and $h\in\operatorname{dom}{B}$, $\langle Bh,A^{*}k\rangle=\langle h,B^{*}A^{*}k\rangle.$ Thus $\langle ABh,k\rangle=\langle h,B^{*}A^{*}k\rangle$. Hence $h\mapsto\langle ABh,k\rangle$ is bounded and $(AB)^{*}k=B^{*}A^{*}k$, proving part (i). If $B^{*}A^{*}\in\mathcal{B}(\mathcal{H})$, then $\operatorname{dom}{(AB)^{*}}\subset\mathcal{H}=\operatorname{dom}{B^{*}A^{*}}$, showing part (ii). ∎
Definition A.20 ([8, Definitions X.2.1 and X.2.3]).
Let $\mathcal{H}$ be a separable Hilbert space. A densely defined operator $A:\mathcal{H}\rightarrow\mathcal{H}$ is said to be symmetric if $\langle Ah,k\rangle=\langle h,Ak\rangle$ for all $h,k\in\operatorname{dom}{A}$. If $A=A^{*}$, then $A$ is said to be self-adjoint.
@@
Lemma A.22 ([8, Proposition X.2.4]).
Let $H$ be a separable Hilbert space and $A$ be a symmetric operator on $\mathcal{H}$.
@@
Let $\mathcal{H}$ be a separable Hilbert space and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{1}(\mathcal{H})_{\mathbb{R}}$ be nonnegative. If $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}\subset\mathcal{H}$ densely, then the following hold.
@@
By Lemma A.22(i), $\mathcal{C}_{1}^{1/2}$ and hence $\mathcal{C}_{1}$ are injective, so (i) holds. By Lemma A.22(ii), $\mathcal{C}_{1}^{-1/2}$ and $\mathcal{C}_{1}^{-1}$ are bijective and self-adjoint. The inverse of a compact operator on its range in an infinite-dimensional space is unbounded, hence (ii) holds. ∎
Condition (i) of the Feldman–Hajek theorem, Theorem 3.1, can be stated equivalently as follows.
@@
Let $\mathcal{H}$ be a Hilbert space and $\mathcal{C}_{1},\mathcal{C}_{2}\in\mathcal{B}(\mathcal{H})$ injective. Then $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$ if and only if $\mathcal{C}_{2}^{-1/2}\mathcal{C}_{1}^{1/2}$ is a well-defined invertible operator in $\mathcal{B}(\mathcal{H})$.
@@
Suppose that $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Then $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is well-defined and bijective. By Lemma A.16(iii), $\mathcal{C}_{1}^{-1/2}$ closed, being the inverse of a bounded, hence closed, operator. By Lemma A.16(i), $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is closed, and by Lemma A.17, it is bounded. Conversely, if $\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}\in\mathcal{B}(H)$ is invertible, then,
$\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ =$=\{\mathcal{C}_{1}^{1/2}h:h\in\mathcal{H}\}=\{\mathcal{C}_{1}^{1/2}\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}h:h\in\mathcal{H}\}=\{\mathcal{C}_{2}^{1/2}h:h\in\mathcal{H}\}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}.$
@@ B.1 Proofs of Section 3
By the Feldman–Hajek theorem, Theorem 3.1, $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Thus, by Lemma A.24, $A\coloneqq\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}$ is a well-defined bounded and invertible operator, and by Theorem 3.1(iii), $AA^{*}-I$ is Hilbert–Schmidt. That is, there exists a sequence $(\lambda_{i})_{i}\subset\ell^{2}(\mathbb{R})$ and ONB $(w_{i})_{i}$ of $\mathcal{H}$ such that,
@@
As $A$ is invertible in $\mathcal{B}(H)$, so are $A^{*}$ and $AA^{*}$. Furthermore, $AA^{*}\geq 0$, hence $AA^{*}>0$ by Lemma A.4, which shows that $\lambda_{i}>-1$ for all $i$, which proves item (i) holds.
@@
Notice that $\frac{-\lambda_{i}}{1+\lambda_{i}}\in\ell^{2}((-1,\infty))$, since $1+\lambda_{i}\rightarrow 1$ and $(\lambda_{i})_{i}\in\ell^{2}((-1,\infty))$. This proves items (ii), (iii) and (iv).
Finally, we prove the statements about the domains of the leftmost operators in items (i), (ii), (iii) and (iv). By Lemma A.23(ii), $\mathcal{C}_{1}^{-1/2}$ is self-adjoint. By Lemma A.19(i), $A^{*}\supset\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}$ and the latter operator is defined on $\operatorname{dom}{\mathcal{C}_{1}^{-1/2}}=\operatorname{ran}{\mathcal{C}_{1}^{1/2}}$ by the definition of composition of linear operators, c.f. Section 1.5. This shows that the leftmost operator in item (i), and by symmetry also in item (iii), is defined on the dense subspace $\operatorname{ran}{\mathcal{C}_{1}^{1/2}}=\operatorname{ran}{\mathcal{C}_{2}^{1/2}}$. Since $A$ is boundedly invertible, $\overline{A^{-1}(\mathcal{D})}=A^{-1}(\overline{\mathcal{D}})=\mathcal{H}$ for any dense set $\mathcal{D}\subset\mathcal{H}$. This shows that $A^{-1}(\operatorname{dom}{\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}})$ is dense in $\mathcal{H}$. Since $\operatorname{dom}{\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}\mathcal{C}_{1}^{-1/2}\mathcal{C}_{2}^{1/2}}=A^{-1}(\operatorname{dom}{\mathcal{C}_{2}^{1/2}\mathcal{C}_{1}^{-1/2}})$, this proves that the leftmost operator in item (ii), and by symmetry also in item (iv), is densely defined. ∎
See 3.7
Proof of Proposition 3.7.
By Lemma 3.4 with $\mathcal{C}_{1}\leftarrow\mathcal{C}_{\textup{pr}}$ and $\mathcal{C}_{2}\leftarrow\mathcal{C}_{\textup{pos}}$, there exists an eigenvalue sequence $(\lambda_{i})_{i\in\mathbb{N}}\subset\ell^{2}((-1,\infty))$ and ONBs $(w_{i})_{i\in\mathbb{N}}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ and items (i), (ii), (iii) and (iv) of Lemma 3.4 hold. In particular, by item (i) and the definition of $R(\cdot\|\cdot)$ in (9), $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})=\sum_{i}\lambda_{i}w_{i}\otimes w_{i}$. By (3c), it holds on $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$,
@@
$\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{C}_{\textup{pr}}^{-1}\mathcal{C}_{\textup{pos}}^{1/2}-I=\mathcal{C}_{\textup{pos}}^{1/2}(\mathcal{C}_{\textup{pr}}^{-1}-\mathcal{C}_{\textup{pos}}^{-1})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(-H)\mathcal{C}_{\textup{pos}}^{1/2},$